OpenStudy (anonymous):

I've been trying to work this problem. i just can't figure it out. "the base of an isosceles triangle is half as long as the two equal sides. write the area of the triangle as a function of the length of the base"

OpenStudy (danjs):



OpenStudy (danjs):

You recall the area is A = 1/2 * base * height draw this line angle bisector ant a right angle forms with the base

OpenStudy (danjs):



OpenStudy (danjs):

that all appear straight forward so far

OpenStudy (anonymous):

ah, now i understand.

OpenStudy (anonymous):

mann, that was easier than i thought. thanks so much.

OpenStudy (phi):

although because they want the area in terms of the base, I would label the base x and a the two equal sides 2x.

OpenStudy (danjs):

yeah i was gonna say that, but figured wait till the end.. . seems like you got it though

OpenStudy (phi):

using ½ * base * height you get x/2 * h you need to find h "in terms of " x

OpenStudy (danjs):

as is , the function will have (x/2) terms,if you change the original base to x and the sides 2x, the area function will have x terms, the overall value is still the same, h will be a bit diff from the pythagoraen calculation

OpenStudy (phi):

yes, but if you label the base x/2 the answer will be in terms of the other side i.e. the expression won't be the one they want.

OpenStudy (danjs):

less you write the final function as \[\sqrt{15}\frac{ x^2 }{ 4 } ---> \sqrt{15}*(\frac{ x }{ 2 })^2\] right A(x/2), i know it is not as nice

OpenStudy (danjs):

maybe not, i see, still defined as half the side, side in terms of the base x is better